I need some help finding the equation of the inverse function if the function is invertible:
f(x)=3(cube root of 9x)
could someone help me with this please?

Answer by stanbon(48535)

I need some help finding the equation of the inverse function if the function is invertible:
f(x)=3(cube root of 9x)
-----------
It is invertible as it is one-to-one.
y=3(9x)^(1/3)
Interchange x and y and solve for y as follows:
x/3 =(9y)^(1/3)
Cube both sides to get:
(x/3)^3 = 9y
y=(1/9)(x/3)^3
y=(1/243)x^3
That is the inverse function.
Cheers,
Stan H.